### v1

# Allocation Multiplicity: Evaluating the Promises of the Rashomon Set

Abstract.
The Rashomon set of equally-good models promises less discriminatory algorithms, reduced outcome homogenization, and fairer decisions through model ensembles or reconciliation. However, we argue from the perspective of allocation multiplicity that these promises may remain unfulfilled. When there are more qualified candidates than resources available, many different allocations of scarce resources can achieve the same utility. This space of equal-utility allocations may not be faithfully reflected by the Rashomon set, as we show in a case study of healthcare allocations. We attribute these unfulfilled promises to several factors: limitations in empirical methods for sampling from the Rashomon set, the standard practice of deterministically selecting individuals with the lowest risk, and structural biases that cause all equally-good models to view some qualified individuals as inherently risky.
allocation multiplicity, model multiplicity, predictive multiplicity, Rashomon set, homogenization, less discriminatory algorithms, algorithmic fairness, underspecification
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06copyright: noneccs: Computing methodologies Philosophical/theoretical foundations of artificial intelligence
1. Introduction
Many decision-making settings involve allocating a fixed number of resources or opportunities among individuals. A company might have hundreds of job applicants but only ten job openings to fill, or a healthcare program may be able to enroll only 100 patients out of thousands. In these cases, there are often more qualified candidates than resources or goods available. This creates allocation multiplicity: the existence of many equal-utility selections of individuals. For example, there are over 3.2 million possible ways to choose 10 qualified people from 25 qualified applicants. The number of possible equal-utility allocations increases even more if we assume that decision-makers will make mistakes and choose some unqualified individuals.
Decision-makers often rely on AI models to determine allocations based on individual risk. The standard practice is to select individuals with the lowest risk or highest likelihood of being qualified based on measurable features and available training data. But as several works have pointed out (Black et al., 2022; Marx et al., 2020; Watson-Daniels et al., 2023b), individuals may receive different risk scores from models with similar training performance, a phenomenon known as model multiplicity. The full space of models with near-equal performance is known as the Rashomon set (Breiman, 2001; Fisher et al., 2019), and a growing literature seeks to find as many models from this set as possible (Watson-Daniels et al., 2023a; Rudin et al., 2024). Using the Rashomon set to inform decision-making may have several benefits. One is the chance to find less discriminatory algorithms (Black et al., 2024b; Gillis et al., 2024) that improve group fairness without sacrificing accuracy. Another is its potential to reduce outcome homogenization (Bommasani et al., 2022; Toups et al., 2023), or the systemic rejection of individuals in multi-shot contexts, if decision-makers adopt different models within the set (Creel and Hellman, 2022; Jain et al., 2024b, a). A third is the opportunity to use the Rashomon set to ensemble (Black and Leino, 2021; Cooper et al., 2024) or reconcile (Du et al., 2024; Roth et al., 2023) conflicting predictions.
Figure 1. Model Multiplicity vs Allocation Multiplicity. Suppose that 10 individuals with different qualifications compete for 5 scarce resources. 2 models from the Rashomon set yield different predictions or risk scores, leading to 2 different equal-utility allocations in which 4 out of 5 selected individuals are qualified. In this setting, there are 60 such equal-utility allocations based on (${6\choose 4}\cdot{4\choose 1}$.
Figure 2. Rashomon Set of Models vs Equal-Utility Allocations. The Rashomon set of all equally-good models is not enumerable, but empirical methods can sample from it. This sample of Rashomon models will lead to some recovered allocations, which may not faithfully represent the full space of equal-utility allocations due to (1) train-test differences, (2) deterministic selection of the lowest risk individuals, and (3) structural biases that cause all near-optimal models to view some qualified individuals as inherently risky.
In this work, we argue that the Rashomon set primarily achieves these aims to the extent that it supports allocation multiplicity. Figure 1 shows that model multiplicity results in different individual-level predictions, which lead to different equal-utility allocations. However, the Rashomon set of models will not yield the full space of equal-utility allocations for several reasons. First, it is unfeasible to enumerate all possible models in the Rashomon set for unbounded hypothesis classes, so we must rely on empirical methods to sample from it (Li et al., 2024; Semenova et al., 2022). This sample of Rashomon models is typically characterized by similar training or validation performance (Watson-Daniels et al., 2023a; Marx et al., 2020), which differs from deployment utility in a resource-constrained setting. Second, decision-makers usually allocate resources by deterministically selecting individuals with the lowest risk, overlooking stochastic mappings between predictions and decisions (Jain et al., 2024a). Third, structural biases in the measurable features and available training data may cause all models in the Rashomon set to consider some qualified individuals as inherently risky. As Figure 2 illustrates, these reasons may lead to recovered allocations from a sample of Rashomon models that are highly unrepresentative of the full space of equal-utility allocations.
The concept of allocation multiplicity enables us to evaluate how well the Rashomon set fulfills its promises of less discriminatory algorithms, reduced outcome homogenization, and fairer decisions through model ensembles or reconciliation. Through a case study of healthcare allocations, we explore various methods for sampling from the Rashomon set and compare their recovered allocations to the full space of equal-utility allocations. Our results demonstrate that: (1) many less discriminatory allocations are not found, (2) outcome homogenization persists in recovered allocations, and (3) the allocation from an ensemble model is unrepresentative of the full space of equal-utility allocations. These findings underscore that allocation multiplicity should serve as a new benchmark for model multiplicity and that recovering equal-utility allocations is essential to fulfilling the promises of the Rashomon set.
2. Model Multiplicity and the Rashomon Set
A growing literature discusses the phenomenon of model multiplicity: the existence of many equally-good models that differ in their individual-level predictions or other properties (Black et al., 2022; Marx et al., 2020; Watson-Daniels et al., 2023b). Equally-good models are models with similar performance for a chosen metric (e.g. accuracy or loss) on a fixed dataset (Ganesh, 2024; Watson-Daniels et al., 2023a). Model multiplicity can arise due to procedural multiplicity, which refers to different choices along the model-building pipeline (Black et al., 2022, 2023). This includes decisions about the problem formulation, feature selection, training data, and model class (Watson-Daniels et al., 2023a; Meyer et al., 2023). However, even when these choices are fixed, model multiplicity can still arise due to uncertainty or noise in the data as well as randomness in the model training process. For example, the data may have been generated through noisy or nondeterministic processes, such as human decision-making (Rudin et al., 2024; Semenova et al., 2023). The data may contain insufficient information for a model to learn generalizable domain knowledge, a problem known as underspecification (D’Amour et al., 2022). Additionally, randomness in the model training process, such as weight initialization, and random seeds, further contribute to model multiplicity (Ganesh et al., 2023).
The Rashomon set [^1] Breiman (2001) named the Rashomon set after the movie “Rashomon,” directed by Akira Kurosawa, in which four different perspectives on the same murder fail to converge to a single truth. consists of all models with near-optimal performance (Breiman, 2001; Fisher et al., 2019; Rudin et al., 2024). While this set generally includes all “good” models, the set can also be defined for a specific hypothesis class, evaluation dataset, and loss function (Hsu and Calmon, 2022; Semenova et al., 2022; Watson-Daniels et al., 2023a). This allows for analyses of its size through metrics like the Rashomon ratio (Semenova et al., 2022) and Rashomon capacity (Hsu and Calmon, 2022). Several empirical methods exist to iterate over the Rashomon set for simple hypothesis classes or sample from it for more complex model types. For example, the Rashomon set is enumerable for decision-trees with a low regularized risk value (Xin et al., 2022), sparse generalized additive models (Zhong et al., 2023), and sparse scoring systems (Liu et al., 2022). Complex hypothesis classes like neural networks have methods to explicitly search for Rashomon models, such as adversarial weight perturbation (Hsu and Calmon, 2022) and dropout-based exploration (Hsu et al., 2024). Other approaches may implicitly produce Rashomon models, such as training on bootstrap datasets (Cooper et al., 2024) or changing random seeds (Ganesh et al., 2023).
2.1. Promises of the Rashomon Set
Previous works discuss several potential benefits of model multiplicity and the Rashomon set (Black et al., 2022; Rudin et al., 2024). Below, we highlight the benefits that are most relevant [^2] Other potential benefits of the Rashomon set include the existence of “simple-yet-accurate” models for interpretability and generalizability (Rudin et al., 2024; Black et al., 2022; Semenova et al., 2022). to our discussion of allocation multiplicity. We refer to these potential benefits as the “promises” of the Rashomon set, as they motivate research on the Rashomon set but may not be fulfilled in practice.
Less Discriminatory Algorithms
Model multiplicity implies that for a given model and a fairness metric, there may exist “less discriminatory algorithms” (LDAs) that are equally-accurate (Black et al., 2022, 2024b). Several works have empirically shown that LDAs exist for fixed populations (Auerbach et al., 2024; Gillis et al., 2024; Coston et al., 2021), but the chosen LDA may be more discriminatory in unseen deployment settings (Black et al., 2024a; Laufer et al., 2024). The goal of searching for LDAs is usually to find a model that improves on group fairness metrics like statistical parity or disparities in error rates without compromising accuracy. This motivates using the Rashomon set to find LDAs (Black et al., 2022; Rudin et al., 2024). In domains such as credit, employment, and housing, decision-makers may have a duty to search for LDAs based on U.S. anti-discrimination laws (Black et al., 2024b). Under the disparate impact doctrine, defendants can justify using a model with discriminatory effects by claiming the model serves a “business necessity,” which often relates to its performance. But even with this justification, defendants may still face liability if the plaintiffs can show there exists a “less discriminatory alternative” that serves the same purpose. In some cases, the courts have placed the burden on defendants to conduct a “reasonable search” for LDAs (Black et al., 2024b; Laufer et al., 2024).
Reduced Outcome Homogenization
Prior works have argued that model multiplicity could help mitigate homogenization across multiple decision-makers (Creel and Hellman, 2022; Jain et al., 2024b, a). Outcome homogenization occurs when the same individuals are subject to consistent errors or negative outcomes (Bommasani et al., 2022; Toups et al., 2023). This represents a grave moral concern because of the possibility that individuals are systemically excluded from important opportunities (Creel and Hellman, 2022; Jain et al., 2024b). Outcome homogenization often arises from algorithmic monoculture, which occurs when decision-makers rely on identical or highly similar models (Kleinberg and Raghavan, 2021; Bommasani et al., 2022). For example, job applicants may receive correlated hiring decisions across companies using similar resume-screening tools. Since the Rashomon set offers several equally-good models, it could reduce homogenization if decision-makers choose models that lead to sufficiently different outcomes for individuals (Creel and Hellman, 2022; Jain et al., 2024b, a). To this end, Creel and Hellman (2022) suggest that decision-makers should randomly choose a model from the Rashomon set. Jain et al. (2024a) propose another approach to randomization that uses equally-good models to first quantify variance in predictions and then randomizes decisions for individuals with high variance.
Fairer Decisions With Ensembles or Reconciliation
Several works use the Rashomon set to ensemble or reconcile predictions (Black and Leino, 2021; Cooper et al., 2024; Roth et al., 2023; Long et al., 2024; Jiang et al., 2023; Du et al., 2024). A common perspective in these works is that choosing and using a single model from the Rashomon set is “arbitrary” [^3] We disagree with this definition of arbitrariness and follow the definition by Creel and Hellman (2022), which considers a model to be arbitrary if it doesn’t serve the purposes of the decision-maker. By this definition, a model from the Rashomon set cannot be arbitrary. However, we agree that using a single model may be unfair for other reasons such as its repeated use at scale (Creel and Hellman, 2022) or its violation of individual claims (Jain et al., 2024a). or unfair to individuals that could have received a more positive prediction under another Rashomon model. To address these concerns, one approach is to ensemble predictions across the Rashomon set of models (Black and Leino, 2021; Cooper et al., 2024; Long et al., 2024; Jiang et al., 2023). Cooper et al. (2024) propose ensembles that abstain from making predictions for individuals with high inconsistencies. Another approach is to reconcile predictions  (Du et al., 2024; Roth et al., 2023) through a process similar to multi-calibration (Hébert-Johnson et al., 2018). Reconciliation considers two models that disagree on individual predictions and uses the disagreement region to iteratively produce a model with better empirical risk minimization.
2.2. Related Work
We build on several related works that evaluate the Rashomon set in deployment settings (Li et al., 2024; Laufer et al., 2024; Watson-Daniels et al., 2023a; Jain et al., 2024a; Simson et al., 2024). Our work differs in that we focus on resource-constrained allocations. Li et al. (2024) distinguish between the theoretical Rashomon set and the sampled Rashomon set from empirical methods in practice. In the context of LDAs, Laufer et al. (2024) highlight how near-optimal models that are less discriminatory for fixed populations may not be less discriminatory for unseen populations. They also characterize the difficulty of finding the least discriminatory algorithm. Jain et al. (2024a) study a similar setting of scarce resource allocations, but focus on the need for randomization in this setting and do not explicitly evaluate models from the Rashomon set. Most similar to our work is that of Watson-Daniels et al. (2023a), who explore model multiplicity under resource constraints. They define the concept of top-$k$ ambiguity, which refers to whether any model from the Rashomon set would place an individual among the top $k$ predicted values. We extend their results beyond linear models and consider top-$k$ ambiguity in the context of all possible equal-utility allocations. This space of possible decisions is similar to what Simson et al. (2024) call multiverse analysis, although their focus is on procedural multiplicity and not a resource-constrained setting.
3. Defining Allocation Multiplicity
In this section, we define the concept of allocation multiplicity and distinguish it from both model multiplicity and the Rashomon set. We also define and quantify the set of all possible equal-utility allocations.
3.1. Preliminaries and Problem Formulation
Consider a setting in which there are $n$ individuals and $k$ resources or opportunities to allocate among them. For example, a company might only have $k$ job openings or a healthcare program might only be able to enroll $k$ patients. An allocation involves the assignment of outcomes $o_{i}\in\{0,1\}$ to each individual $i$. Importantly, there is scarcity in that not all individuals can receive positive outcomes with $k\ll n$.
Definition 0.
An allocation $\mathbf{o}=(o_{1},o_{2},\dots,o_{n})$ involves the assignment of outcomes $o_{i}\in\{0,1\}$ to each individual $i$ such that only $k$ out of $n$ individuals receive positive outcomes, under a selection rate $\frac{k}{n}$. The space of all allocations is:
$\mathcal{O}_{k,n}=\left\{\mathbf{o}\in\{0,1\}^{n}\,|\,\sum^{n}_{i=1}o_{i}=k\right\}\,\text{where}\,\,|\mathcal{O}_{k,n}|={n\choose k}$
Individuals may have varying qualifications $o^{*}_{i}$ for the resources or opportunities being allocated. We simplify to the case where each individual is qualified ($o^{*}_{i}=1$) or unqualified ($o^{*}_{i}=0$). For example, a patient may be qualified for enrollment in a healthcare program if they have a chronic illness. In many settings, the number of qualified individuals $n^{\prime}$ is greater than the number of available resources or opportunities $k$.
Definition 0.
The qualification rate $\frac{n^{\prime}}{n}$ is the number of qualified individuals $n^{\prime}$ out of the total number of candidate individuals $n$.
Decision-makers do not know whether individuals are qualified at the time of allocation. For many decision-makers, the objective [^4] When $k$ is fixed, our formulation of utility is equivalent to other formulations (Gillis et al., 2024; Laufer et al., 2024) that assign a benefit for selecting qualified individuals and a penalty for selecting unqualified individuals. However, we do not consider the possibility of varying benefits across qualified individuals. is to maximize the number of qualified individuals selected in an allocation, which we denote using $k^{\prime}$.
Definition 0.
The utility or precision of an allocation is $\frac{k^{\prime}}{k}=\frac{1}{k}\sum_{i=1}^{n}o^{*}_{i}\cdot o_{i}$, which is simply the number of selected qualified individuals $k^{\prime}$ out of the total number of selected individuals $k$.
Risk-averse decision-makers aim to maximize utility by selecting individuals with the highest probability of being qualified. While the true probability $p_{i}=\mathbb{P}(o^{*}_{i}=1)$ is unknowable [^5] Individual probabilities misrepresent qualifications because the events in question are typically realized only once (Dawid, 2017; Dwork et al., 2021; Roth et al., 2023)., the conditional probability $p(x_{i})=\mathbb{P}(o^{*}_{i}=1\,|\,x_{i})$ can be estimated using data. The validity of taking $p(x_{i})$ to be an estimate of $p_{i}$ depends on the choice of features and how well qualifications can be predicted from them. Setting these specifications aside, machine learning models can estimate $p(x_{i})$ based on features $x_{i}$ from a given domain $\mathcal{X}$.
Definition 0.
A model $\mathbf{\hat{p}}:\mathcal{X}\rightarrow[0,1]$ maps an individual’s features $x_{i}\in\mathcal{X}$ to a prediction $\hat{p}(x_{i})$, which estimates the conditional probability $p(x_{i})=\mathbb{P}(o^{*}_{i}=1\,|\,x_{i})$.
In data-driven allocations, outcomes $o_{i}$ are assigned based on predictions $\hat{p}(x_{i})$. Standard practice in machine learning is to deterministically assign $o_{i}=1$ to individuals with the top $k$ predictions. However, there are various other ways to map predictions to outcomes. For example, a human-in-the-loop could determine outcomes for high uncertainty predictions (Cooper et al., 2024), or a weighted lottery based on predictions could determine the outcomes (Jain et al., 2024a).
Definition 0.
A mapping =$\mathcal{M}=\{(\hat{p}(x_{1}),o_{1}),(\hat{p}(x_{2}),o_{2}),\dots,(\hat{p}(x_{n}),o_{n})\}$ characterizes the relationship between a model $\mathbf{\hat{p}}$ and allocation $\mathbf{o}$. In a top $k$ mapping, $o_{i}=1$ if and only if $\,\hat{p}(x_{i})$ is in the $k$ highest predictions.
3.2. From Model Multiplicity to Allocation Multiplicity
Model multiplicity refers to the existence of multiple models with similar performance but different properties, including producing different predictions for certain individuals (a subtype of model multiplicity called “predictive multiplicity”)  (Marx et al., 2020; Black et al., 2022; Watson-Daniels et al., 2023b). Performance is measured on a training or validation dataset $\mathcal{D_{\texttt{train}}}=\{(x_{i},o^{*}_{i})\}^{n_{\texttt{train}}}_{i=1}$ because the target variable $o^{*}_{i}$ (in our case, qualification) is not known at the time of allocation. Following definitions of model multiplicity in prior work (Marx et al., 2020; Watson-Daniels et al., 2023a), we consider a baseline model $\mathbf{\hat{p}_{0}}$ that is the solution to an empirical risk minimization problem of the form $\min_{\mathbf{\hat{p}}\in\mathcal{H}}L(\mathbf{\hat{p}};\mathcal{D_{\texttt{train}}})$, over a hypothesis class $\mathcal{H}$ with loss function $L(\cdot\,;\mathcal{D})$. For example, $\mathcal{H}$ could be the set of neural networks represented by a specific architecture and $L(\cdot\,;\mathcal{D})$ could be the cross-entropy loss. In this context, the $\epsilon-$Rashomon set encompasses all models that achieve near-optimal loss.
Definition 0.
For a baseline model $\mathbf{\hat{p}_{0}}$ and loss tolerance $\epsilon>0$, the $\mathbf{\epsilon}$-Rashomon set of models is:
$\mathcal{R}_{\epsilon}(\mathbf{\hat{p}_{0}}):=\{\mathbf{\hat{p}}\in\mathcal{H}:\mathbb{L}(\mathbf{\hat{p}};\mathcal{D_{\texttt{train}}})\leq\mathbb{L}(\mathbf{\hat{p}_{0}};\mathcal{D_{\texttt{train}}})+\epsilon\}$
and model multiplicity exists if $\,|\mathcal{R}_{\epsilon}(\mathbf{\hat{p}_{0}})|>1$. An $\mathbf{\epsilon}$-Rashomon model is a model $\mathbf{\hat{p}}\in\mathcal{R}_{\epsilon}(\mathbf{\hat{p}_{0}})$.
Building on the concept of model multiplicity, we define allocation multiplicity as the existence of multiple allocations with near-equal deployment utility. Recall that an allocation $\mathbf{o}$ is the assignment of outcomes to individuals (Definition 3.1) and that its utility depends on $k^{\prime}$: the number of selected individuals who are qualified (Definition 3.3). Let $\mathbf{o_{0}}$ denote an allocation based on a mapping $\mathcal{M}_{0}$ from the baseline model’s predictions $\mathbf{\hat{p}_{0}}$. In this context, the set of $\Delta$-equal-utility allocations encompasses all allocations that achieve near-equal utility.
Definition 0.
For a baseline allocation $\mathbf{o_{0}}$ and utility tolerance $\Delta\geq 0$, the set of $\Delta$-equal-utility allocations is:
$\mathcal{R}_{\Delta}(\mathbf{o_{0}}):=\left\{\mathbf{o_{\delta}}\in\mathcal{O}_{k,n}\,|\,k^{\prime}_{\delta}\geq k^{\prime}_{0}-\Delta\right\}$
and allocation multiplicity exists if $\,|\mathcal{R}_{\Delta}(\mathbf{o_{0}})|>1$. A $\mathbf{\Delta}$-equal-utility allocation is an allocation $\mathbf{o}\in\mathcal{R}_{\Delta}(\mathbf{o_{0}})$. When $\Delta=0$, $\mathcal{R}_{0}(\mathbf{o_{0}})$ is a set of equal-utility allocations.
An $\epsilon$-Rashomon model does not guarantee a $\Delta$-equal-utility allocation for several reasons. First, $\epsilon$ measures deviation from the optimal training or validation loss, while $\Delta$ measures deviation from a specific deployment utility. These metrics are not directly comparable given that the loss is not evaluated in a resource-constrained setting (Watson-Daniels et al., 2023a; Jain et al., 2024a). Generalization error adds further uncertainty, as training or validation performance might differ from deployment utility. Moreover, predictions can be mapped to outcomes in various ways, such as through stochastic procedures or with a human-in-the-loop. Thus, even a model outside the $\epsilon$-Rashomon set could yield a $\Delta$-equal-utility allocation. For example, a human decision-maker might have domain knowledge for high-uncertainty predictions, helping them decide on a high-utility allocation when using a sub-optimal model. Even for the top $k$ mapping, a model could make predictions with high loss (e.g. predictions clustered around 0.5), yet maintain an ordering that yields high utility.
3.3. Quantifying Allocation Multiplicity
The number of equal-utility allocations is often exponentially high. We can calculate how many equal-utility allocations are possible based on the number of total individuals $n$, qualified individuals $n^{\prime}$, resources $k$, and qualified individuals selected $k^{\prime}$ under the baseline allocation $\mathbf{o}_{0}$. An allocation will have the same utility as $\mathbf{o}_{0}$ if it selects exactly $k^{\prime}$ qualified individuals and $k-k^{\prime}$ unqualified individuals. This represents a combinatorial problem, as there are $n^{\prime}$ qualified individuals to choose $k^{\prime}$ from, and $n-n^{\prime}$ unqualified individuals to choose $k-k^{\prime}$ from.
Proposition 3.8.
For a setting in which there are $n$ individuals, $k$ resources, $n^{\prime}$ qualified individuals, and $k^{\prime}$ qualified individuals selected by a baseline allocation $\mathbf{o}_{0}$, the number of $\mathbf{\Delta}$-equal-utility allocations is:
$|\mathcal{R}_{\Delta}(\mathbf{o}_{0})|=\sum_{k^{\prime}_{\delta}\,=\,k^{\prime}-\Delta}^{k^{\prime}}{n^{\prime}\choose k^{\prime}_{\delta}}\cdot{n-n^{\prime}\choose k-k^{\prime}_{\delta}}$
Number of Equal-Utility Allocations $|\mathcal{R}_{0}(\mathbf{o}_{0})|$
$k^{\prime}/k$ $n^{\prime}/n$ $n=100$ $n=1000$
$k/n=0.10$ $k/n=0.25$ $k/n=0.50$ $k/n=0.10$ $k/n=0.25$ $k/n=0.50$
[…]
Table 1. $|\mathcal{R}_{0}(\mathbf{o}_{0})|$ for a baseline allocation $o_{0}$ with varying utility $k^{\prime}/k$, qualification rate $n^{\prime}/n$, and selection rate $k/n$.
[…]
4. Why Does Allocation Multiplicity Matter?
Now that we have seen how many equal-utility allocations are possible, we discuss why allocation multiplicity matters. First, risk-averse decision-makers might assume that all or most good equal-utility allocations will be produced by the set of equally-good models. Allocation multiplicity highlights that “riskier” allocations can be chosen without sacrificing utility. Second, decision-makers who prioritize normative ideals other than utility may value allocation multiplicity for reasons aligned with their principles. Third, those who view the Rashomon set positively for its potential benefits may find that the space of all equal-utility allocations is what actually delivers those benefits.
4.1. From the Perspective of Risk-Aversion
Risk-averse decision-makers will usually choose the allocation with the highest expected utility based on a model’s predictions. And from the perspective of the information they have at the time, this might be the most utility-maximizing step they can take. But model multiplicity means that even in the set of models that are built with the same problem formulation and trained on the same data, there are many models that could minimize risk to an equivalent extent, thereby delivering equivalent expected utility to the risk-averse. We take the next step, pointing out that realized utility, not the expected utility, is what ultimately matters to most decision-makers. While our current information does not always allow us to find all the equal-utility allocations, the knowledge that they exist should spur decision-makers to collect the additional information needed to recover more, and a more even sampling, of the equal-utility allocations. Comparing each highest-expected-utility model with the set of all possible equal-utility allocations in the training data is a tool to highlight the shortcomings of the current training data and the inherent limitations of allocating based on solely on minimizing risk.
From the Perspective of Claims
Allocation multiplicity has a positive role in ensuring that the use of machine learning models to make decisions lives up to our normative commitments. For example, we might be morally committed to fulfilling claims that individuals have to goods or outcomes for reasons of fairness (Broome, 1990). According to the claims view, an individual has a claim to an organ if she needs it more than others; a claim to a job if she merits it; or a claim to a good if she was promised it and therefore deserves it (Broome, 1990). If resources are scarce, not all claims can be fulfilled. Our comparison of model multiplicity to allocation multiplicity show that drawing predictions only from the Rashomon set models over-respects the claims of those who appear less risky and under-respects the claims of those who appear riskier. But those with equal claims deserve equal chances at the good. Although we may not always have the information needed to predict allocations correctly, the claims framework suggests that we should (1) seek out additional information that would allow us to better respect all claims and (2) be prepared to give “compensatory” goods to those whose claims we failed to respect.
From the Perspective of the Worst-Off
Another variant of utilitarianism that can guide us in creating and choosing allocations is prioritarianism. Prioritarianism holds that when allocating resources, we should prioritize improvements to the well-being of the worst-off (Arneson, 2022). Considering equal-utility allocations allows prioritiarians to help the worst-off in two ways. First, within the set of apparently equal-utility allocations, there are likely allocations that give more resources to those who suffer the most. Thus just as the search for the LDA can be used to increase fairness by choosing the fairest of the equally-good models, the prioritarian may search the equal-utility allocation space for an allocation that prioritizes the worst-off. Second, the shift in perspective from model multiplicity to allocation multiplicity suggests that many of the “multiple” models have allocations that are concentrated on the least-risky individuals. Those individuals are likely to be the best-off, not the worst off. So prioritarianism would suggest using the allocation-multiplicity perspective to spur the search for the seemingly-risky people who would be most benefited by an allocation.
(a) Less Discriminatory Algorithms
(b) Outcome Homogenization
Figure 3. Evaluating Promises of the Rashomon Set from the Perspective of Allocation Multiplicity: (a) the least discriminatory equal-utility allocation may not be recovered; (b) outcome homogenization may persist in recovered allocations.
4.2. From the Perspective of the Rashomon Set
Prior works suggest that the Rashomon set promises benefits such as less discriminatory algorithms, reduced outcome homogenization, and fairer decisions through ensembles or reconciliation. However, we contend that allocations derived from the Rashomon set (e.g., via a top-$k$ mapping) may fall short of fully delivering on these promises. To the extent that the Rashomon set helps us deliver these benefits, it is because it helps us to recover more allocations from the broader space of equal-utility allocations. In this section, we argue that considering equal-utility allocations directly can better achieve these aims.
Less Discriminatory Algorithms
For a given fairness metric, there exists a least discriminatory allocation within the space of all equal-utility allocations. For example, decision-makers may accidentally under-select individuals who are racial minorities because they are perceived as riskier by predictive models. However, excluding these individuals is not necessary: provided there are at least $k^{\prime}$ qualified individuals within the minority population, an equal-utility allocation exists that consists entirely of minority individuals. But since the all, majority, or even proportional minority allocations are difficult for predictive models to find, this may cause all “less discriminatory algorithms” (LDAs) in the Rashomon set to result in allocations that deviate significantly from the “least discriminatory” equal-utility allocation, as seen in Figure 3(a). Moreover, generalization error further contributes to this issue: prior research has shown that LDAs derived from training or validation data can become more discriminatory when deployed in unseen settings (Black et al., 2024a; Laufer et al., 2024).
Reduced Outcome Homogenization
In the space of all equal-utility allocations, every individual has a non-zero probability of being selected when allocations do not achieve perfect utility. This means that the systemic rejection rate could theoretically be reduced to zero without sacrificing utility, provided there are enough decision-makers. However, as illustrated in Figure 3(b), allocations derived from the Rashomon set may still exclude some qualified individuals entirely, leaving them with zero chance of selection. This underscores that predictive multiplicity, as captured by the Rashomon set, does not fully deliver on its promise of reducing homogenization if this is a primary concern. Viewing the problem through the lens of allocation multiplicity makes this limitation clear and highlights how considering the space of equal-utility allocations can provide better insight into and potential pathways for reducing outcome homogenization.
Fairer Decisions With Ensembles or Reconciliation
Ensembling or reconciling predictions across models in the Rashomon set is often proposed as a way to achieve fairer decisions. However, these methods are limited when viewed through the lens of allocation multiplicity. While they can or smooth predictions across models, they may fail to capture the full range of equal-utility allocations that the Rashomon set implicitly supports. Ensembling or reconciling predictions necessarily is based on the starting set of Rashomon predictions, but the Rashomon set is not an even representation of the set of equal-utility allocations. A default decision-making approach—such as using a top-$k$ mapping may not account for the diversity of allocations or represent the average characteristics of individuals chosen across the space of equal-utility allocations. This limitation underscores the need to consider allocation multiplicity directly, as it provides a more comprehensive view of fairness by encompassing all plausible and representative outcomes within the space of equal-utility allocations.
5. Case Study: Healthcare Allocations
We turn to a case study of healthcare allocations in order to analyze allocation multiplicity in a real-world setting. Specifically, we evaluate how well the Rashomon set fulfills its promises from the perspective of allocation multiplicity. Our results show that: (1) many less discriminatory allocations are not found, (2) outcome homogenization persists in recovered allocations, and (3) the allocation from an ensemble model is unrepresentative of the full space of equal-utility allocations.
5.1. Data and Methods
We use a healthcare dataset [^6] Obermeyer et al. (2019) compiles data for all primary care patients at a large academic hospital. Due to the sensitivity of this data, the study authors released a semi-synthetic version that is designed to closely mirror the original dataset. This data is available at: https://gitlab.com/labsysmed/dissecting-bias. released by Obermeyer et al. (2019) that was used to select patients for a “high-risk care management” program at a hospital. The hospital aimed to identify high-risk patients in a given year based on information about the patient’s health and healthcare costs in the previous year. Specifically, the dataset contains features [^7] In our experiments, we use the same subset of features as Watson-Daniels et al. (2023b) (see Appendix Table 5). In particular, we do not use race as a feature. for demographics (age, race, sex), chronic illnesses that a patient had in the previous year, and costs claimed by the patients’ insurer in the previous year. These costs are broken down for various healthcare services, such as primary care, emergency room visits, and pharmacy expenses.
Different Bootstrapped Shuffled Weight
Feature Subsets Training Data Data Order Perturbation
$\frac{n^{\prime}}{n}$ $\frac{k}{n}=0.25$ $\frac{k}{n}=0.50$ $\frac{k}{n}=0.25$ $\frac{k}{n}=0.50$ $\frac{k}{n}=0.25$ $\frac{k}{n}=0.50$ $\frac{k}{n}=0.25$ $\frac{k}{n}=0.50$
Number of Rashomon Models 0.32 1674 1674 997 997 959 959 999 999
0.55 1386 1386 1000 1000 835 835 1000 1000
Validation Loss 0.32 0.279 0.279 0.284 0.284 0.284 0.284 0.290 0.290
0.55 0.327 0.327 0.328 0.328 0.330 0.330 0.335 0.335
Number of Recovered Allocations 0.32 502 373 997 997 75 62 54 68
0.55 734 460 1000 1000 184 74 26 31
Allocation Utility ($k^{\prime}/k$) 0.32 0.873 0.602 0.877 0.604 0.877 0.605 0.877 0.604
0.55 0.950 0.902 0.953 0.904 0.952 0.905 0.952 0.905
Table 2. Performance Metrics For Different Methods of Sampling From the Rashomon Set. All methods produce many models with near-equal performance, but some methods yield more unique allocations (based on a top $k$ mapping) than others.
Using this data, we run various simulations of a scenario in which a hospital is allocating enrollment in a healthcare program among its patients. In order to determine who is qualified for the program (Obermeyer et al., 2019; Watson-Daniels et al., 2023a), the hospital could predict which patients will have the highest future healthcare costs or the most chronic illnesses. However, Obermeyer et al. (2019) demonstrate that using predicted healthcare costs to select patients can result in racial bias. We therefore consider an individual to be qualified if they will have $q$ or more chronic illnesses in the upcoming year. To test different qualification rates, we vary $q$ to represent having at least 1, 2, or 3 chronic illnesses, which corresponds to 55%, 32%, and 19% of the population, respectively. Not all qualified patients can be selected because, as in the real-world [^8] In the real-world hospital studied by Obermeyer et al. (2019), patients with risk scores in the top 3% were automatically enrolled in the program, while those in the top 45% were flagged for their doctors, who could enroll them based on program availability., healthcare programs are resource-constrained (Obermeyer et al., 2019; Watson-Daniels et al., 2023a). To simulate this, we consider different selection rates of 10%, 25%, and 50%. For each combination of qualification and selection rates, we run 250 simulations [^9] We use 10 different partitions of a 60-20-20 train-validation-test split. For each partition, we run 25 simulations where a different set of $n=1000$ patients are drawn from the test set to be candidates for our hypothetical allocation. with different test sets of $n=1000$ patients as allocation candidates.
In each simulation, we consider the Rashomon set of models that predict whether individuals are qualified. Since this set is non-enumerable, we evaluate various empirical methods for sampling from the Rashomon set. We choose these methods because they represent different points along the ML pipeline that may lead to model multiplicity (Black et al., 2023, 2022).
  • Different Feature Subsets: We follow the procedure in Liu et al. (2022) to explore different sparse scoring systems (FasterRisk). These are linear models with integer coefficients and a restricted number of features.
  • Bootstrapped Training Data: We follow the procedure in Cooper et al. (2024) to train different neural networks based on bootstrapped sub-samples of the training data.
  • Shuffled Data Order: We follow the procedure in Ganesh et al. (2023) to shuffle the data order between neural network training epochs, and take the predictions at each epoch to be a different “model.” This is an efficient way to approximate prediction variance from different weight initializations (Ganesh et al., 2023).
  • Weight Perturbation: We follow the procedure in Hsu and Calmon (2022) to adversarially perturb neural network weights. Perturbed models are obtained by fine-tuning to increase the predicted score for different validation data points until the validation loss exceeds $\epsilon$.
For each method, we consider the best-performing model in terms of cross-entropy loss on a validation dataset. We define the empirical sample of Rashomon models to be all models with a validation loss within $\epsilon=0.01$ of the best-performing model for a given method [^10] The theoretical Rashomon set can vary across methods, as it is defined based on the best-performing model for that method. We choose this approach because the Rashomon set is often defined for a specific hypothesis class (Watson-Daniels et al., 2023a; Marx et al., 2020). However, in our simulations, all the methods yield models with an average validation loss around 0.01 of each other, as Table 2 shows.. We assume the recovered allocations from this sample are determined using a top $k$ mapping, where individuals with the highest $k$ predictions are selected. For each method, Table 2 shows the number of models found in the Rashomon set and the corresponding number of recovered allocations. Note that many models (unique sets of predictions) map to the same allocation based on the top $k$ mapping. Table 2 also shows the average utility of recovered allocations, which is the proportion of selected patients with $q$ or more chronic illnesses. This average utility serves as the basis for defining our reference of all equal-utility allocations.
(a) The “least discriminatory” equal-utility allocation is not found based on the threshold test, which evaluates if the ratio of chronic illnesses between selected black and white patients is greater than 1.
(b) Outcome homogenization persists in recovered allocations from the Rashomon set: between 15% to 30% of qualified individuals are systemically rejected.
(c) The age of selected patients is skewed older in the average allocation from the Rashomon set, indicating how ensembling predictions may lead to unrepresentative allocations.
Figure 4. Comparing Recovered Allocations From the Rashomon Set to All Equal-Utility Allocations ($\frac{k}{n}=0.25$, $\frac{n^{\prime}}{n}=0.32$)
$k/n$ $n^{\prime}/n$ All Equal-Utility Different Bootstrapped Shuffled Weight
Allocations Feature Subsets Training Data Data Order Perturbation
Least Discriminatory Allocation 0.25 0.32 0.86 1.12 1.12 1.19 1.18
0.25 0.55 0.92 1.10 1.12 1.18 1.20
0.50 0.32 1.21 1.26 1.25 1.31 1.28
0.50 0.55 0.90 1.25 1.24 1.30 1.29
Outcome Homogenization 0.25 0.32 0.80 0.95 0.97 0.99 0.99
0.25 0.55 0.71 0.92 0.95 0.98 0.99
0.50 0.32 0.69 0.95 0.96 0.99 0.98
0.50 0.55 0.76 0.94 0.96 0.99 0.99
Age Homogenization 0.25 0.32 2.49 2.35 2.33 2.36 2.36
0.25 0.55 2.56 2.20 2.21 2.31 2.33
0.50 0.32 2.52 2.51 2.48 2.47 2.47
0.50 0.55 2.55 2.49 2.46 2.45 2.45
Table 3. Evaluating the Promises of the Rashomon Set. The “least discriminatory” allocation is based on the threshold test metric, which evaluates the ratio of chronic illnesses between selected black and white patients. Outcome homogenization measures the pairwise consistency of outcomes for individuals. Age homogenization measures the entropy over the age distribution of selected patients, where a lower entropy indicates a more narrow and unrepresentative distribution.
5.2. Evaluating Promises of the Rashomon Set
For each Rashomon set sampling method, we compare its allocations to the full set of equal-utility allocations, revealing that the recovered allocations fall short of the Rashomon set’s potential.
The least discriminatory allocation from the Rashomon set is far from the least discriminatory allocation with equal-utility that could be achieved.
In real-world healthcare allocations, Obermeyer et al. (2019) highlight the fact that selected black patients are considerably sicker than selected white patients in terms of how many active chronic illnesses they have. This indicates discrimination based on the threshold test, which evaluates if the ratio of active chronic illnesses between selected black and white patients is greater than 1. In addition to its unfairness, this allocation is concerning from a prioritarian point of view: there are many black patients who were not selected who are sicker than the white patients who were selected. Across recovered allocations from the Rashomon set, we consider the least discriminatory allocation to have the smallest ratio based on the threshold test. Figure 4(a) shows that the smallest ratio recovered by all 4 methods is around 1.2, indicating that this “least discriminatory allocation” is still discriminatory. In contrast, the space of all equal-utility allocations contains many allocations with threshold test ratios at or less than 1. We specifically compare to an ideal equal-utility allocation that first selects $k^{\prime}$ qualified patients in descending order by number of chronic illnesses, then selects $k-k^{\prime}$ unqualified patients in the same fashion, alternating between black and white patients when there are ties. As Figure 4(a) shows, this results in a least [^11] We consider this the “least discriminatory” equal-utility allocation because it is the least discriminatory according to the normative goal of selecting patients based on how sick they are. Strictly speaking, an equal-utility allocation with a lower ratio could be obtained by adversarially choosing the least-ill black patients and the most-ill white patients. discriminatory equal-utility allocation with a threshold test ratio around 0.9.
Even if decision-makers use different models from the Rashomon set, outcome homogenization remains high and some qualified individuals are still systemically rejected.
We compare the pairwise consistency of individual outcomes between recovered allocations from the Rashomon set and all equal-utility allocations. Pairwise consistency represents the probability an individual would receive the same outcome under two random allocations from the set under consideration (Cooper et al., 2024). As Table 3 shows, the pairwise consistency is around 95% for recovered allocations from the Rashomon set. This is much higher than the pairwise consistency across all equal-utility allocations, which is around 75%. If decision-makers exclusively choose models from the Rashomon set, the high pairwise consistency implies that many individuals may be systemically rejected, even if they are qualified. Figure 4(b) analyzes the selection probability for qualified individuals across recovered allocations from the Rashomon set. Depending on the method, between 15% to 30% of qualified individuals are never selected in any allocation from the Rashomon set.
The ensemble allocation from the Rashomon set is unrepresentative of all equal-utility allocations.
Consider the scenario where a fairness-oriented decision-maker uses the Rashomon set to average predictions and then selects patients with the top $k$ highest predictions from this ensemble. We show that the selected individuals under this allocation have different features from the selected individuals under a randomly-drawn equal-utility allocation. We focus on a specific feature – the age of selected patients – because of its potential to be homogenized in healthcare allocations. A stark example is the UK’s liver transplant matching algorithm, which systematically denied transplants to younger patients regardless of medical urgency (Narayanan et al., 2024). Figure 4(c) shows a similar result for the ensemble allocation from the Rashomon set: the age of selected patients is skewed older. While older patients are expected to be sicker and more qualified, Figure 4(c) also shows the age distribution of selected patients under a randomly-drawn equal-utility allocation, which chooses significantly more younger patients. Table 3 measures this phenomenon using the Shannon entropy of age brackets across selected patients. The recovered allocations from the Rashomon set have a lower entropy, indicating more homogenization in the age of selected patients.
6. Discussion
(a) Many qualified individuals have “high-risk” or low prediction values, indicating structural biases or underspecification in their risk scores.
(b) The lowest-risk allocation and the 10,000-th lowest-risk allocation, along with the allocations in between, have utility within 0.01 of each other on average (shaded region is S.D.).
Figure 5. (a) Risk Distribution of Qualified Individuals; (b) Low Deviations in Utility for Many “Riskier” Allocations
Our results show that conventional methods of using the Rashomon set do not fulfill its promises in resource-constrained allocations. Since we have established that model multiplicity is only useful insofar as it yields allocation multiplicity, how can we find more equal-utility allocations? We discuss three approaches below as avenues for future research.
Improving Sampling from the Rashomon Set
Allocation multiplicity should serve as a new benchmark for empirical methods that aim to find Rashomon models, especially given that a growing area of work involves developing and improving such methods (Rudin et al., 2024; Black et al., 2022). As Table 2 shows, all the methods that we explore find many different models in the Rashomon set. However, some methods perform better than others in terms of how many unique allocations they recover under a top $k$ mapping. We find that using different subsets of features and training data recovers more unique allocations than shuffling data order or perturbing model weights. But given that the former methods involve re-training different models from scratch, there are computational limitations to their use in practice. Even when many unique allocations are recovered, we further show that they fall short of what the full-space of equal-utility allocations promises. We hope the perspective of allocation multiplicity inspires future work to develop methods of finding models that deliver on these promises.
Mitigating Structural Biases and Underspecification
Addressing structural biases and underspecification in how risk scores are generated may lead to equal-utility allocations that status-quo models fail to reflect. Even if we sample extensively from the Rashomon set, the features used to model risk may lack the richness needed to distinguish between high-risk and low-risk individuals, particularly within certain subgroups. For example, in healthcare allocation, models trained on datasets that underrepresent low-income populations may fail to account for systemic barriers, such as limited access to preventive care or higher exposure to environmental risks (Obermeyer et al., 2019). As a result, these individuals are often misclassified as high-risk, even when they might benefit greatly from treatment (c.f. Figure 5(a)). This deficiency is compounded by incomplete or biased training data that limits the ability of models to learn meaningful patterns for marginalized groups. Consequently, even diverse models from the Rashomon set can converge on overly simplistic or biased risk estimates, systematically excluding certain qualified individuals and narrowing the recovered allocation space.
Non-Deterministic Mappings Between Predictions and Decisions
Our case study focused on the limitations of allocations recovered by deterministically selecting individuals with the lowest risk. However, there are many stochastic procedures to map predictions to decisions (Jain et al., 2024a). Decision-makers could use a weighted lottery based on predicted risk scores, or randomize decisions for individuals near the top $k$ threshold. These methods may better explore the space of allocations than trying to find more models in the Rashomon set, which are not guaranteed to produce different allocations under the deterministic paradigm. While stochastic procedures may involve trade-offs with expected utility, there may be little or no trade-off with realized utility, especially given the exponential number of equal-utility allocations that exist (c.f. Figure 5(b)). We hope that the knowledge of allocation multiplicity inspires a paradigm shift in how decision-makers interpret and act on model predictions, emphasizing flexibility and equity over rigid, risk-based thresholds.